OpenStudy (anonymous):

If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its radius is....

OpenStudy (amistre64):

x^2+y^2 = r^2 .... compare

OpenStudy (anonymous):

no its the formula for the circle with its centre at its origin

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

he is right but a small change (x-h)^+(y-k)^=r^ r^=36 r=6

OpenStudy (amistre64):

a circles radius doent change just becuase you center it someplace else .....

OpenStudy (anonymous):

that's correct the euation you gave can be used to find the radius but if we expand and find something else it would be wrong